# A Unified Implicit Attention Formulation for Gated-Linear Recurrent Sequence Models

### v1

@@ 1 Introduction
The very recent State Space Model (SSM) named Mamba by Gu and Dao [20] has attracted considerable attention since its recent debut [31, 32, 64, 30, 43, 60, 40, 63], further establishing it as an efficient and accurate general-purpose model. Like other SSM models [21, 24, 22], Mamba is autoregressive during inference and trains efficiently through convolutions. Recently, Ali et al. [2] have highlighted a third aspect of the Mamba model; namely, that it is also an attention model, since it implicitly computes attention.
@@
Let $X$ be the matrix whose columns are the hidden values associated with each token, and let $\alpha$ be the matrix of mixing coefficients. The set of values of the next layer is initially obtained as $Y=\alpha X$ and it can then undergo other forms of processing, such as nonlinear activations and per-token processing.
@@
Given a neural architecture, one always linearizes mixing operators and writes them in the form $Y=\alpha X$. However, to be considered an attention model it is required that $\alpha$ be a function of $X$, which means that the linear operator is data-dependent. This property is shown by Ali et al. [2] to hold only for the recent selective SSM (S6) [20], but not for earlier SSMs. Specifically, for standard state-space layers, it has been demonstrated that they can be linearized into a constant operator, represented by a constant matrix $alpha$, which is solely controlled by the layer’s parameters. However, in the S6 layers, $alpha$ is influenced by both the input and the layer’s parameters.
@@
The implicit attention matrix of Ali et al. [2] considers the S6 mechanism and ignores the influence of other critical mixer components, such as Conv1D, gate branch, linear layers, and SiLU activations. The formulation we propose in this work incorporates these additional elements and, as we show empirically, leads to improved interpretability results in both computer vision and NLP.
@@
Furthermore, using a similar holistic formulation, we show that S6 is not the only sequence model that implicitly computes attention and that an implicit attention representation can also describe other recent layers, such as RWKV [41], Griffin [14] ,HGRN [47] and more, as illustrated in Fig. 1
@@
Our main contributions are as follows: (i) We introduce the implicit self-attention representation, unifying transformers with non-transformer layers, such as Griffin [14], RWKV [41], ReNet [52], and others. (ii) We refine the approach of [2] to produce more accurate attention matrices. The previous work focused exclusively on the S6 layer, without considering the gating and Conv1D sub-layers in Mamba, while our representation incorporates all these factors. (iii) While “Attention is not Explanation” [27], transformer explainability relies heavily on attention matrices. We demonstrate that our implicit attention representation of non-transformer models can be used to develop new explainability and interpretability techniques for non-transformer models, enhancing the community’s ability to understand, explore, and manage aspects of robustness, bias, fairness, and safety. As a sample downstream application, we demonstrate excellent out-of-the-box results for weakly supervised semantic segmentation. (iv) Finally, our framework facilitates comparisons between transformers and other recent architectures, by providing a unified attention view and setting the stage for further improvements and insights.
@@ 2 Related Work
Self-Attention
Self-attention, a cornerstone of Transformer architectures [55], has profoundly influenced recent developments in NLP and computer vision. This mechanism leverages pairwise token interactions to dynamically allocate focus across different parts of the input sequence, assessing the relevance of each token in relation to others. The computational formula is given by:
@@
Applications of Attention Matrices
Attention matrices play a crucial role in transformers, as multiplying these matrices with value vectors is the core operation that captures interactions between tokens. Beyond this essential role in computing self-attention, they are also used for various purposes: (i) Explainability and Interpretability: Although attention itself is not inherently explainable [27], many methods in these domains rely on attention matrices to understand and analyze model behavior [1, 11, 10, 2] . (ii) Multi-modal Learning: Numerous multi-modal learning schemes are based on variations of cross-attention, enabling dependencies to be learned between any pair of tokens of different modalities [33, 53]. (iii) Weakly Supervised Tasks: Attention matrices can provide a valuable source of supervision, highlighting relevant regions or relationships within the data to guide model learning. These techniques are popular in semantic segmentation [48, 59, 49], and robustness enhancement [12]. Finally, (iv) Inductive Bias and Regularization Methods: Since attention matrices represent interactions between tokens, they inherently carry semantic meaning. Therefore, they can be manipulated to incorporate domain knowledge or regulate the model effectively [29, 4, 9, 65].
S6 Layers and Mamba
The recently presented selective SSM [20] (S6) outperforms the previous SSMs and various other architectures in NLP [3, 58], vision [32, 64], graph classification [56, 6], and more. S6 incorporates a dynamic input-dependent form of the discrete matrices $\bar{A},\bar{B},$ and $C$, such that for every time-step the SSM employs a different recurrent rule. This technique differs from the previous state-space layers, which use the same set of matrices and recurrent rules for each time step.
@@
where $S_{B},S_{C},S_{\Delta}$ are linear projection layers, and SoftPlus is the smooth elementwise ReLU approximation. While other SSMs employ complex values and non-diagonal matrices, S6 employs a real-diagonal parametrization.
The usage of input-dependent time-variant layers adds to the expressivity of the layer, allowing it to adapt to the input, and potentially captures more complex dependencies. While other input-dependent time-variant mechanisms have been proposed in previous works through gated RNNs, the S5 layer [50], or adaptive filtering via input-dependent IIR filters [34], S6 also presents an efficient IO-aware implementation, which is parallelized on GPUs via work-efficient parallel scanners [8, 36].
@@
$\hat{x}=\text{SiLU( Conv1D( Linear(}x\text{) ) )},\quad\hat{z}=\text{SiLU( Linear(}x\text{) )},\quad\hat{y}^{\prime}=\text{Linear}(\text{Selective SSM}(\hat{x})\otimes\hat{z})),\quad$ (3)
@@
The vision mamba architectures [32, 64] (ViM) follow the vision transformer (ViT) [15] but replace the Transformer’s self-attention mechanism by two Mamba layers, where each layer is applied in a bidirectional manner. These vision models outperform the standard ViT model in terms of accuracy and efficiency, for models of similar parameter counts.
Gated-Linear RNNs
RNNs, along with their advanced versions, such as GRU [13] and LSTM [26], play a fundamental role in deep sequence modeling. Their auto-regressive design decouples sequence length from computational complexity per step, making them highly efficient at decoding. However, they don’t scale as effectively as transformers and often face challenges, such as slow training and vanishing gradients. Recently, linear RNNs have shown improved abilities in capturing long-range dependencies [21, 39] and enhanced scalability [42, 14]. Furthermore, gated linear RNNs deliver surprisingly strong language modeling performance [37, 57, 41, 47]. The most advanced gated linear RNNs include the following variants: (i) RWKV-6 [41], which draws inspiration from attention-free transformers (AFT) [62], (ii) Mamba [20], which employs selective SSM, (iii) HGRN2 [46], which utilizes state expansion, and (iv) Hawk [14], which is built upon an enhanced variant of the LRU [39]. Other notable examples include GLA [61], GateLoop [28], and RenNet [52]. These layers achieve results comparable to transformers on larger scales, matching well-known models, such as Pythia [7] and LLaMA 2 [54]. Moreover, several studies show that hybrid models combining attention mechanisms with gated linear RNNs can be complementary [14, 31, 45, 35, 5, 19], enhancing both approaches. Despite these successes, interoperability and explainability techniques for these models remain relatively unexplored.
@@ 3 Method
In this section, we present a general and holistic data-control linear operator representation that can be applied to (at least) many of the recent non-transformer architectures and which incorporates all components of the architecture. Sec. 3.1 formulates the entire Mamba architecture as a data-control linear operator, incorporating subcomponents such as Conv1D, gate branches, and activations. Subsequently, Sections. 3.2 and 3.3 extend our approach to other architectures, such as Griffin [14] and RWKV [41]. Additionally, in Appendix A, we present how to extract holistic data-controlled attention matrices for RetNet [52] and HGRN [47].
@@ 3.1 Formulation of Mamba via Attention matrices
Mamba can be formulated in a way that separates the components that mix channels from those that mix tokens, see Fig. 3 in [20].
$\textbf{Mamba}(x)=\textbf{Linaer}_{3}\Big{(}\textbf{ SILU}(\textbf{Linear}_{2}(\textbf{Linear}_{1}(x)))\otimes\textbf{S6}(\textbf{SILU}(\textbf{Conv1D}(\textbf{Linear}_{1}(x))))\Big{)}$ (4)
Since $Linear_{1}$ and $Linear_{3}$ do not mix tokens, they are irrelevant (similarly to the MLP layers) and we consider the simplified expression:
$\textbf{Mamba}(x)=\Big{(}\textbf{ SILU}(\textbf{Linear}_{2}(x))\Big{)}\otimes\Big{(}\textbf{S6}(\textbf{SILU}(\textbf{Conv1D}(x)))\Big{)}$ (5)
@@
$\textbf{Mamba}(x)=\textbf{diag}\Big{(}\textbf{ SILU}(\textbf{Linear}_{2}(x))\Big{)}\Big{(}\textbf{S6}(\textbf{SILU}(\textbf{Conv1D}(x)))\Big{)}$ (6)
The S6 layer can be formalized as a data-control linear operator (see Eq. 12 in [2]):
@@
By plugging Eq. 7 into Eq. 6 and since $\textbf{SILU}(x)=\textbf{Sig}(x)\cdot x$:
$\textbf{Mamba}(x)=\underbrace{\textbf{diag}\Big{(}\textbf{ SILU}(\textbf{Linear}_{2}(x))\Big{)}}_{W_{x}^{\prime}\in\mathbb{R}^{L\times L},\quad(\textbf{gate)}}\hat{\alpha}\underbrace{\textbf{diag}\Big{(}\textbf{Sig}(\textbf{Conv1D}(x))\Big{)})}_{Z_{x^{\prime}}\in\mathbb{R}^{L\times L},\quad(\textbf{Conv \&Act})}(\textbf{Conv1D}(x))$ (8)
@@
$\textbf{Mamba}(x)={W_{x}^{\prime}}{\hat{\alpha}}{Z_{x^{\prime}}}Mx=Hx,\quad H={W_{x}^{\prime}}{\hat{\alpha}}{Z_{x^{\prime}}}M$ (9)
@@ 3.2 Formulation of Griffin via Attention Matrices
$y=\textbf{Linear}_{3}\Big{(}\Big{(}\textbf{GeLU}(\textbf{Linear}_{1}(x^{\prime}))\Big{)}\otimes\Big{(}\textbf{RG-LRU}(\textbf{Conv1D}(\textbf{Linear}_{2}(x^{\prime}))\Big{)}\Big{)}$ (10)
@@
$x=\textbf{Linear}_{2}(x^{\prime}),\quad y=\textbf{Linear}_{3}\Big{(}\textbf{diag}\Big{(}\textbf{GeLU}(\textbf{Linear}_{1}^{\prime}(x)\Big{)}\Big{(}\textbf{RG-LRU}(\textbf{Conv1D}(x)\Big{)}\Big{)}$ (11)
Note that $\textbf{Linear}_{1}^{\prime}:=\textbf{Linear}_{1}\textbf{Linear}_{2}$ and $\textbf{Linear}_{3}$ do not mix tokens and can therefore be omitted. By substituting Conv1D with matrix multiplication using a causal convolution matrix $M$, we derive:
$y=\textbf{diag}\Big{(}\textbf{GeLU}(\textbf{Linear}_{1}^{\prime}(x)\Big{)}\Big{(}\textbf{RG-LRU}(Mx)\Big{)}$ (12)
@@
$h=\tilde{\alpha}x,\quad\begin{bmatrix}h_{1}\\ h_{2}\\ \vdots\\ h_{L}\\ \end{bmatrix}=\begin{bmatrix}\sqrt{1-{a_{1}}^{2}}\otimes i_{1}&0&\cdots&0\\ a_{2}\sqrt{1-{a_{1}}^{2}}\otimes i_{1}&\sqrt{1-{a_{2}}^{2}}\otimes i_{2}&\cdots&0\\ \vdots&\vdots&\ddots&0\\ \Pi_{k=2}^{L}a_{k}\sqrt{1-{a_{1}}^{2}}\otimes i_{1}\quad&\Pi_{k=3}^{L}a_{k}\sqrt{1-{a_{2}}^{2}}\otimes i_{2}\quad&\cdots\quad&\sqrt{1-{a_{L}}^{2}}\otimes i_{L}\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{L}\\ \end{bmatrix}$ (14)
@@
$y=\textbf{diag}\Big{(}\textbf{GeLU}(\textbf{Linear}_{1}^{\prime}(x)\Big{)}\tilde{\alpha}Mx=Hx,\quad H=\textbf{diag}\Big{(}\textbf{GeLU}(\textbf{Linear}_{1}^{\prime}(x)\Big{)}\tilde{\alpha}M$ (15)

### v2

# Explaining Modern Gated-Linear RNNs via A Unified Implicit Attention Formulation

@@ 1 Introduction
The very recent State Space Model (SSM) named Mamba by Gu & Dao (2023) has attracted considerable attention since its recent debut (Lieber et al., 2024; Liu et al., 2024; Zhu et al., 2024; Xu et al., 2024), further establishing it as an efficient and accurate general-purpose model. Like other SSM models (Gu et al., 2021a; b), Mamba is autoregressive during inference and trains efficiently in parallel. Recently, Ali et al. (2024) have highlighted a third aspect of the Mamba model; namely, that it is also an attention model, since it implicitly computes attention.
@@
Let $X$ be the matrix whose columns are the hidden values associated with each token, and let $\alpha$ be the matrix of mixing coefficients. The set of values of the next layer is initially obtained as $Y=\alpha X$ and it can then undergo other forms of processing, such as nonlinear activations and per-token processing. Given a neural architecture, one can always linearize the mixing operators and write them in the form $Y=\alpha X$ via their first-order approximation. However, to be considered an attention model it is required that $\alpha$ be a function of $X$, which means that the linear operator is data-dependent. This property is shown by Ali et al. (2024) to hold only for the recent selective SSM (S6) , but not for most earlier SSMs. Specifically, for standard state-space layers, it has been demonstrated that they can be linearized into a constant operator, represented by a constant matrix $\alpha$, which is solely controlled by the layer’s parameters. However, in the S6 layers, $\alpha$ is influenced by both the input and the layer’s parameters.
@@
The implicit attention matrix of Ali et al. (2024) considers the S6 mechanism and ignores the influence of other critical mixer components, such as Conv1D, gate branch, linear layers, and SiLU activations. The formulation we propose in this work incorporates these additional elements and, as we show empirically, leads to improved interpretability results in both computer vision and NLP.
@@
Furthermore, using a similar holistic formulation, we show that S6 is not the only sequence model that implicitly computes attention and that an implicit attention representation can also describe other recent layers, such as RWKV (Peng et al., 2023), Griffin (De et al., 2024) ,HGRN (Qin et al., 2024b) and more, as illustrated in Figure 1.
To achieve a more accurate representation that better reflects the model’s behavior, we employ a composition of multiple components. The concept of composing non-attention layers and representing them as data-controlled linear operators was initially introduced in Hyena (Poli et al., 2023), which attempts to replicate attention capabilities through a composition of two sub-quadratic operators (long convolutions and multiplicative gating). Our formulation differs from this approach in two main ways. First, instead of focusing on replicating attention capabilities, we take the reverse step by demonstrating that, through a sequence of algebraic manipulations, several existing modern gated linear RNNs can be viewed as single implicit attention layers. Second, our goal is to find the most accurate implicit attention representation possible, as it crucial for applications like interpretability. This leads to a significant extension over Hyena’s work. For example, while Hyena’s matrices are constructed from the two components mentioned above, our implicit attention representation incorporates additional layers, including non-linear operators such as linear layers, activations, short convolutions, and normalization layers. For instance, our formulation for Mamba-2 is built upon six different layers, some of which appear multiple times, resulting in a much more complex outcome. Additionally, while other works explore the relations between non-attention layers and linear attention (Arora et al., 2023), we not aware to any work extending the concept of composition of components similar to us, or apply it to existing modern RNN such as Griffin, Mamba, or link in to interpateblity or similar domains.
@@
Our main contributions are as follows: (i) We introduce the implicit self-attention representation, unifying Transformers with non-Transformer layers, such as Griffin, RWKV, ReNet, and others. (ii) We refine the approach of Ali et al. (2024) to produce more accurate attention matrices. The previous work focused exclusively on the S6 layer, without considering the gating and Conv1D sub-layers in Mamba, while our representation incorporates all these factors (and additional peripherals in other models) (iii) While “Attention is not Explanation” (Jain & Wallace, 2019), Transformer explainability relies heavily on attention matrices. We demonstrate that our implicit attention representation of non-Transformer models can be used to develop new explainability and interpretability techniques for non-Transformer models, enhancing the community’s ability to understand, explore, and manage aspects of robustness, bias, fairness, and safety. As a sample downstream application, we demonstrate excellent out-of-the-box results for attribution-based performance-enhancing techniques. (iv) Finally, our framework facilitates comparisons between Transformers and other recent architectures, by providing a unified attention view and setting the stage for further improvements and insights.
@@ 2 Related Work
Self-Attention.  Self-attention, a cornerstone of Transformer architectures (Vaswani, 2017), has profoundly influenced recent developments in NLP and computer vision. This mechanism leverages pairwise token interactions to dynamically allocate focus across different parts of the input sequence, assessing the relevance of each token in relation to others. The computational formula is given by:
@@
Applications of Attention Matrices.  Attention matrices play a crucial role in Transformers, as multiplying these matrices with value vectors is the core operation that captures interactions between tokens. Beyond this essential role in computing self-attention, they are also used for various purposes: (i) Explainability and Interpretability: Although attention itself is not inherently explainable (Jain & Wallace, 2019), many methods in these domains rely on attention matrices to understand and analyze model behavior (Abnar & Zuidema, 2020; Chefer et al., 2021b; a; Ali et al., 2024) . (ii) Multi-modal Learning: Numerous multi-modal learning schemes are based on variations of cross-attention, enabling dependencies to be learned between any pair of tokens of different modalities (Lu et al., 2019; Tan & Bansal, 2019). (iii) Weakly Supervised Tasks: Attention matrices can provide a valuable source of supervision, highlighting relevant regions or relationships within the data to guide model learning. These techniques are popular in semantic segmentation (Ru et al., 2022; Wang et al., 2020; Ru et al., 2023), and robustness enhancement (Chefer et al., 2022). Finally, (iv) Inductive Bias and Regularization Methods: Since attention matrices represent interactions between tokens, they inherently carry semantic meaning. Therefore, they can be manipulated to incorporate domain knowledge or regulate the model effectively (Li et al., 2018; Attanasio et al., 2022; Bonaldi et al., 2023; Zimerman & Wolf, 2023).
S6 Layers and Mamba. The recently presented selective SSM (Gu & Dao, 2023) (S6) outperforms the previous SSMs and various other architectures in NLP (Anthony et al., 2024; Wang et al., 2024b), vision (Liu et al., 2024; Zhu et al., 2024), graph classification (Wang et al., 2024a; Behrouz & Hashemi, 2024), and more. S6 incorporates a dynamic input-dependent form of the discrete matrices $\bar{A},\bar{B},$ and $C$, such that for every time-step the SSM employs a different recurrent rule. This technique differs from the previous state-space layers, which use the same set of matrices and recurrent rules for each time step.
@@
where $S_{B},S_{C},S_{\Delta}$ are linear projection layers, and Softplus is asmooth approximation of ReLU.
The usage of input-dependent time-variant layers adds to the expressivity of the layer, allowing it to adapt to the input, and potentially captures more complex dependencies. While other input-dependent time-variant mechanisms have been proposed in previous works through gated RNNs, the S5 layer (Smith et al., 2022), or adaptive filtering via input-dependent IIR filters (Lutati et al., 2023), S6 also presents an efficient IO-aware implementation, which is parallelized on GPUs via work-efficient parallel scanners (Blelloch, 1990; Martin & Cundy, 2017).
@@
$\hat{x}=\text{SiLU( Conv1D( Linear(}x\text{) ) )},\quad\hat{z}=\text{SiLU( Linear(}x\text{) )},\quad\hat{y}^{\prime}=\text{Linear}(\text{Selective SSM}(\hat{x})\otimes\hat{z}),\quad$ (3)
@@
The vision Mamba architectures (Liu et al., 2024; Zhu et al., 2024) (ViM) follow the vision Transformer (ViT) (Dosovitskiy et al., 2020) but replace the Transformer’s self-attention mechanism by two bidirectional Mamba layers, These vision models outperform the standard ViT in terms of accuracy and efficiency, for models of similar parameter counts.
Gated-Linear RNNs. RNNs, along with their advanced versions, such as GRU (Chung et al., 2014) and LSTM (Hochreiter & Schmidhuber, 1997), play a fundamental role in deep sequence modeling. Their auto-regressive design decouples sequence length from computational complexity per step, making them highly efficient at decoding. However, they don’t scale as effectively as Transformers and often face challenges, such as slow training and vanishing gradients. Recently, linear RNNs have shown improved abilities in capturing long-range dependencies (Gu et al., 2021a; Orvieto et al., 2023) and enhanced scalability (Peng et al., 2024; De et al., 2024). Furthermore, gated linear RNNs deliver surprisingly strong language modeling performance (Mehta et al., 2022; Wang et al., 2022; Peng et al., 2023; Qin et al., 2024b). The most advanced gated linear RNNs include the following variants: (i) RWKV-6 (Peng et al., 2023), which draws inspiration from attention-free Transformers (AFT) (Zhai et al., 2021), (ii) Mamba (Gu & Dao, 2023), which employs selective SSM, (iii) HGRN2 (Qin et al., 2024a), which utilizes state expansion, and (iv) Hawk (De et al., 2024), which is built upon an enhanced variant of the LRU (Orvieto et al., 2023). Other notable examples include GLA (Yang et al., 2023), GateLoop (Katsch, 2023), and RenNet (Sun et al., 2023). These layers achieve results comparable to Transformers on larger scales, matching well-known models, such as Pythia (Biderman et al., 2023) and LLaMA 2 (Touvron et al., 2023). Moreover, several studies show that hybrid models combining attention mechanisms with gated linear RNNs can be complementary (De et al., 2024; Lieber et al., 2024; Poli et al., 2024; Ma et al., 2022; Baron et al., 2023; Fu et al., 2022), enhancing both approaches. Despite these successes, interoperability and explainability techniques for these models remain relatively unexplored.
@@ 3 Method
In this section, we present a general and holistic data-control linear operator representation that can be applied to (at least) many of the recent non-Transformer architectures and which incorporates all components of the architecture. Our objective is to describe each layer in the form of $y=\tilde{\alpha}x$ such that x and y is the input and output respectively, and $\tilde{\alpha}=f(x;\Theta_{\text{arch}})$ is an attention matrix controlled by the parameters of the model and the input. Sec. 3.1 formulates the entire Mamba and Mamba-2 Dao & Gu (2024) architectures as a data-control linear operator, incorporating subcomponents such as Conv1D, gate branches, normalizations and activations. Subsequently, Sections. 3.2 and 3.3 extend our approach to other architectures, such as Griffin (De et al., 2024) and RWKV (Peng et al., 2023). Additionally, in Appendix A, we present how to extract holistic data-controlled attention matrices for RetNet (Sun et al., 2023) and HGRN (Qin et al., 2024b).
@@ 3.1 Formulation of Mamba via Attention matrices
Mamba can be formulated in a way that separates the components that mix channels from those that mix tokens:
$\text{Mamba}(x)=\text{Linear}_{3}\Big{(}\text{ SILU}(\text{Linear}_{2}(\text{Linear}_{1}(x)))\otimes\text{S6}(\text{SILU}(\text{Conv1D}(\text{Linear}_{1}(x))))\Big{)}$ (4)
Since $\text{Linear}_{1}$ and $\text{Linear}_{3}$ do not mix tokens, they are less relevant to our representation (similar to the MLP layers in Transformers), and we consider the following simplified expression:
$\text{Mamba}(x)=\Big{(}\text{ SILU}(\text{Linear}_{2}(x))\Big{)}\otimes\Big{(}\text{S6}(\text{SILU}(\text{Conv1D}(x)))\Big{)}$ (5)
@@
$\text{Mamba}(x)=\text{diag}\Big{(}\text{ SILU}(\text{Linear}_{2}(x))\Big{)}\Big{(}\text{S6}(\text{SILU}(\text{Conv1D}(x)))\Big{)}$ (6)
The S6 layer can be formalized as a data-control linear operator (see Eq. 12 in (Ali et al., 2024)):
@@
By plugging Eq. 7 into Eq. 6 and since $\text{SILU}(x)=\text{Sigmoid}(x)\cdot x$:
$\text{Mamba}(x)=\underbrace{\text{diag}\Big{(}\text{ SILU}(\text{Linear}_{2}(x))\Big{)}}_{W_{x}^{\prime}\in\mathbb{R}^{L\times L},\quad(\text{gate)}}\hat{\alpha}\underbrace{\text{diag}\Big{(}\text{Sigmoid}(\text{Conv1D}(x))\Big{)}}_{Z_{x^{\prime}}\in\mathbb{R}^{L\times L},\quad(\text{Conv \& Act})}(\text{Conv1D}(x))$ (8)
@@
$\text{Mamba}(x)={W_{x}^{\prime}}{\hat{\alpha}}{Z_{x^{\prime}}}Mx=Hx,\quad H={W_{x}^{\prime}}{\hat{\alpha}}{Z_{x^{\prime}}}M$ (9)
@@
Mamba-2 This architecture builds upon Mamba by introducing two key enhancements relevant to our formulation: (i) incorporating the concept of multiple heads via a multi-input SSM, and (ii) applying additional normalization (GroupRMSNorm) after the multiplicative gating.
The first modification can be handled by broadcasting parts of the equations across different attention heads. For the second modification, we first compute the per-head statistics necessary for Group Normalization and pack them into a diagonal matrix.
$\mu_{h}=\frac{1}{d}\sum_{i=1}^{d}x_{h}[i],\quad\sigma_{h}=\epsilon+\sqrt{\frac{1}{d}\sum_{i=1}^{d}\left(x_{h}[i]-\mu_{h}\right)^{2}},\quad\mathbb{N}=\text{diag}\left(\frac{1}{\sigma_{1}},\cdots,\frac{1}{\sigma_{h}},\cdots,\frac{1}{\sigma_{H}}\right)$ (10)
where $x_{h}[i]$ denotes the $i$-th feature of head $h\in[H]$, $d$ is the dimensionality of each head, and $\epsilon$ is a small constant added for numerical stability in GroupRMSNorm.
The matrix $\mathbb{N}$ allows us to represent the GroupRMSNorm operator via matrix multiplication such that $\mathbb{N}x=\text{GroupRMSNorm}(x)$ (where $\mathbb{N}$ is augmented across groups). Thus, we plug these modifications into our formulation of Mamba (Equation 9), obtaining the following implicit-attention formulation for Mamba-2:
[…]
@@ 3.2 Formulation of Griffin via Attention Matrices
$y=\text{Linear}_{3}\Big{(}\Big{(}\text{GeLU}(\text{Linear}_{1}(x^{\prime}))\Big{)}\otimes\Big{(}\text{RG-LRU}(\text{Conv1D}(\text{Linear}_{2}(x^{\prime}))\Big{)}\Big{)}$ (12)
@@
$x=\text{Linear}_{2}(x^{\prime}),\quad y=\text{Linear}_{3}\Big{(}\text{diag}\Big{(}\text{GeLU}(\text{Linear}_{1}^{\prime}(x))\Big{)}\Big{(}\text{RG-LRU}(\text{Conv1D}(x))\Big{)}\Big{)}$ (13)
Note that $\text{Linear}_{1}^{\prime}:=\text{Linear}_{1}\text{Linear}_{2}$ and $\text{Linear}_{3}$ do not mix tokens and can therefore be omitted. By substituting Conv1D with matrix multiplication using a causal convolution matrix $M$, we derive:
$y=\text{diag}\Big{(}\text{GeLU}(\text{Linear}_{1}^{\prime}(x)\Big{)}\Big{(}\text{RG-LRU}(Mx)\Big{)}$ (14)
@@
$h=\tilde{\alpha}x,\textbf{ }\begin{bmatrix}h_{1}\\ h_{2}\\ \vdots\\ h_{L}\\ \end{bmatrix}=\begin{bmatrix}\sqrt{1-{a_{1}}^{2}}\otimes i_{1}&0&\cdots&0\\ a_{2}\sqrt{1-{a_{1}}^{2}}\otimes i_{1}&\sqrt{1-{a_{2}}^{2}}\otimes i_{2}&\cdots&0\\ \vdots&\vdots&\ddots&0\\ \Pi_{k=2}^{L}a_{k}\sqrt{1-{a_{1}}^{2}}\otimes i_{1}\quad&\Pi_{k=3}^{L}a_{k}\sqrt{1-{a_{2}}^{2}}\otimes i_{2}\quad&\cdots\quad&\sqrt{1-{a_{L}}^{2}}\otimes i_{L}\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{L}\\ \end{bmatrix}$ (16)
@@
$y=\text{diag}\Big{(}\text{GeLU}(\text{Linear}_{1}^{\prime}(x))\Big{)}\tilde{\alpha}Mx=Hx,\quad H=\text{diag}\Big{(}\text{GeLU}(\text{Linear}_{1}^{\prime}(x))\Big{)}\tilde{\alpha}M$ (17)